### v1

@@ 6.5.1 Training LLMs on Dhoroni
The advent of large language models (LLMs) has transformed the field of natural language processing (Fan et al., 2023), offering powerful tools for understanding and generating human language (Zhou and Chen, 2023). Training LLMs on the Dhoroni dataset presents a significant opportunity to advance the capabilities of NLP for the Bengali language, particularly in the context of climate change and environmental issues. By leveraging the rich data contained within the Dhoroni dataset, researchers can develop LLMs that are fine-tuned to recognize and generate Bengali text related to climate topics.
@@ 6.5.2 Performance Improvement in Low-Performing Models
To improve the poor performance in tasks like Climate/Environmental Topics Detection, News Target Detection, and Authority Involvement Detection, several modeling approaches can be explored. These tasks involve complex patterns that current methods may struggle to capture. Multi-task learning can enhance performance by training related tasks together, allowing models to share learned features. For example, training on News Target Detection and Authority Involvement Detection simultaneously may help the model generalize better. Lastly, further hyperparameter tuning and regularization techniques, such as learning rate schedulers, dropout, and weight decay, can optimize model performance and reduce overfitting. Ensembling multiple models may also improve robustness by leveraging the strengths of different approaches for better overall performance.
@@ 7 Conclusion
Novelty and Relevance Statement
Dhoroni dataset and BanglaBERT-Dhoroni model family represent a pioneering effort in utilizing natural language processing for Bengali climate discourse, addressing a significant research gap in environmental management. By providing helpful insights into local climate narratives, our work enhances understanding and promotes effective communication strategies that are essential for informed decision-making. This initiative not only advances knowledge in the field but also fosters broader engagement with climate issues among diverse audiences.
Conflict of Interest
The authors declare that there are no conflicts of interest regarding the publication of this paper. All research procedures followed ethical guidelines, and the study was conducted with integrity and transparency. The authors have no financial, personal, or other relationships that could inappropriately influence or bias the content of this work.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used ChatGPT in order to reduce grammatical errors and writing clarity. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
@@
Data related to this work is publicly available at DOI: 10.5281/zenodo.13695110 (Wasi et al., 2024) under the MIT Licence.
@@
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. We are grateful to everyone who supported us throughout the data collection, annotation, modeling, and writing processes.
Acknowledgement
The authors are grateful to the Computational Intelligence and Operations Laboratory (CIOL) for their support, as well as for the initial review and feedback on the manuscript.
[…]

### v2

@@ 6.5 Future Work Directions
For future work, we propose two broad areas: utilizing Large Language Models and improving Low-Performing Models.
@@ 6.5.1 Training LLMs on Dhoroni
The advent of large language models (LLMs) has transformed the field of natural language processing (Fan et al., 2023), offering powerful tools for understanding and generating human language (Zhou & Chen, 2023). Training LLMs on the Dhoroni dataset presents a significant opportunity to advance the capabilities of NLP for the Bengali language, particularly in the context of climate change and environmental issues. By leveraging the rich data contained within the Dhoroni dataset, researchers can develop LLMs that are fine-tuned to recognize and generate Bengali text related to climate topics.
@@ 6.5.2 Performance Improvement in Low-Performing Models
To improve poor performance in tasks such as detecting climate / environmental topics, detecting news targets, and detecting involvement of the authority, several modeling approaches can be explored. These tasks involve complex patterns that current methods may struggle to capture. Multi-task learning can enhance performance by training related tasks together, allowing models to share learned features. For example, training in news target detection and authority involvement detection simultaneously may help the model generalize better. Lastly, further hyperparameter tuning and regularization techniques, such as learning rate schedulers, dropout, and weight decay, can optimize model performance and reduce overfitting. Ensembling multiple models may also improve robustness by leveraging the strengths of different approaches for better overall performance.
@@ 7 Conclusion
Acknowledgement
The authors thank the Computational Intelligence and Operations Laboratory (CIOL) for their support, as well as for the initial review and feedback on the manuscript.
Author Contributions
A.T.W. and W.F. contributed to the work equally. A.T.W.: Conceptualization of this study, Methodology, Data curation, Resources, Investigation, Formal analysis, Supervision, Validation, Visualization, Writing – original draft, Writing – review & editing, Project administration. W.F.: Software, Methodology, Formal analysis, Investigation, Resources, Data Curation, Project administration. T.A.: Resources, Data Curation. A.R.: Resources, Data Curation, Writing – original draft, Writing – review & editing. M.R.I.: Resources, Data Curation, Writing – review & editing.
Competing Interests
The authors declare that there are no competing interest regarding the publication of this paper. All research procedures followed ethical guidelines, and the study was conducted with integrity and transparency. The authors have no financial, personal, or other relationships that could inappropriately influence or bias the content of this work.
@@ Acknowledgement
Data related to this work is publicly available in Zenodo at DOI: 10.5281/zenodo.13695110 (Wasi et al., 2024) under the MIT license.
Materials Availability
Trained models are publicly available as Collection in Hugging Face at ID: ciol-research/dhoroni-6723c1c675c6ef21285c0dda under the MIT license.
@@
This work received no specific grant from any funding agency, commercial or not-for-profit sectors.